# Mapping atmospheric features of the planetary-mass brown dwarf SIMP 0136 with JWST NIRISS

### v1

@@ Appendix B Results of the regional spectra retrieval
To avoid uncorrelated noise at individual wavelengths we average the spectra in wavelength bins, then construct maps from lightcurves of binned data. We created maps from 400 light curves of 0.0045 bins using 4th order spherical harmonics. We then select 6 regions of interest and extract the spectra from 400 maps at respective longitude and latitude. The resulting spectra and a map with regions of interest are shown on Fig.10.
We then apply the same atmospheric retrieval framework as described in Section 5 to explore T-P profiles, molecular abundances and cloud levels of each region. As seen from Fig.11 the temperature profiles retrieved from regional points tend to be more isothermal in the deep atmosphere, and fail to reproduce the adiabatic slope seen in the T-P profile from the time-average integrated spectrum retrieval in Section 5. As a result, the altitudes and fractions of forsterite clouds vary significantly for each region as well. We also detect an anti-correlation, where the darkest cloud patch at lat: 0, lon: -60 has the hottest T-P profile (cyan curve in Fig.11), while the hottest point at lat: 0, lon: 15 has the lowest T-P profile (grey curve). Molecular abundances are also very different for spectrum of each region (see Figure 12).

### v2

@@ Appendix A PCA explained variance and BIC
Two independent analyses that were used to identify the number of SIMP 0136 variability sources are shown in Figure 9. The first principal component of the PCA explains 50% of data variance, and the second—31%, while the rest of the components are insignificant. According to Cowan et al. (2011), this means there are 3 spectral regions that influence the observed signal variability. This is supported by the BIC analysis during matching linear combinations of Sonora Diamondback models to our time-averaged spectrum, which suggests that six spectral regions is the most optimal case.
@@ Appendix B Results of the regional spectra retrieval
To avoid uncorrelated noise at individual wavelengths we average the spectra in wavelength bins, then construct maps from lightcurves of binned data. We created maps from 400 lightcurves of 0.0045 bins using fourth-order spherical harmonics. We then select six regions of interest and extract the spectra from 400 maps at the respective longitude and latitude. The resulting spectra and a map with regions of interest are shown on Fig.10.
We then apply the same atmospheric retrieval framework as described in Section V to explore T-P profiles, molecular abundances and cloud levels of each region. As seen from Fig.11 the temperature profiles retrieved from regional points tend to be more isothermal in the deep atmosphere, and fail to reproduce the adiabatic slope seen in the T-P profile from the time-averaged integrated spectrum retrieval in Section V. As a result, the altitudes and fractions of forsterite clouds vary significantly for each region as well. We also detect an anti-correlation, where the darkest cloud patch at latitude: 0, longitude: -60 has the hottest T-P profile (cyan curve in Fig.11), while the hottest point at latitude: 0, longitude: 15 has the lowest T-P profile (grey curve). Molecular abundances are also very different for spectrum of each region (see Figure 12).